A 2.50 kg object falls vertically downward in a viscous medium at a constant speed of 2.50 m/s. How much work is done by the force the viscous medium exerts on the object as it falls 80.0 cm?
Answer #1

Weight Acting Downward = M*g
Force due to viscous medium on object, = Fv

As the object is moving with constant velocity, Acceleration is zero

Fv = M*g
Fv = 2.5 * 9.8 m/s^
Fv = 24.5 N

Now, Work done = Force *Distance
W = 24.5 * 0.8
W = 20.4 J
Work is done by the force the viscous medium exerts on the object, W = 20.4 J